OpenStudy (anonymous):

You and two friends got jobs at stores in the mall. You earn $117.60 for 14 hours of work. Ty earns $148.50 in 18 hours. Kim earns $137.60 in 16 hours. Who has the highest pay rate in dollars per hour? You all earn at the same rate. Kim You Ty

OpenStudy (anonymous):

You : 117.60/14 = $8.4/hour Kim: 137.60/16 = $8.6/hour Ty: 148.50/18 = $8.25/hour